# Certification of high stellar ranks of quantum states of light with a pair of click detectors

### v3

@@ VI Conclusions
In conclusion, we have shown that high stellar rank of certain quantum states of light can be certified with a very simple experimental setup that consists of a beam splitter, two click detectors, and an attenuator. We have explained how the reduced total detection efficiency helps to certify stellar rank of higher Fock states, and we have numerically calculated the stellar rank thresholds for the considered stellar rank witnesses. For the simple witness based on probability that D1 clicks and D2 does not click we have observed that the maximum detectable stellar rank increases with decreasing detection efficiency and increases with increasing beam splitter transmittance. The ability to certify higher stellar ranks can be further improved by considering a more refined stellar rank witnesses that combine probability $R_{1}$ and probability $R_{2}$ that both detectors click simultaneously. In this latter case stellar ranks higher than $1$ can be certified even with perfect detectors and $\eta=1$ by utilizing an unbalanced scheme with $T\neq 0.5$. Nevertheless, even in such case losses can be helpful. Our results provide new insights into the capabilities of the multiplexed click detectors and reveal that high stellar ranks can be certified even with a very simple scheme that involves only two detectors.
@@ Acknowledgements.
The author acknowledges support by the project OP JAC CZ.02.01.01/00/23_021/0008790 of the Ministry of Education, Youth, and Sports of the Czech Republic and EU.
@@
Here we disucss the properties of a parametric curve specified by $R_{2}(n)$ and $R_{1}(n)$, where $n$ is now treated as a continuous parameter and $T=\frac{1}{2}$. Recall that
@@
which is positive for $10<\eta<$ 1. Since $R_{2}(n)$ monotonically increases with $n$, in order to prove the concavity it suffices to consider the signed curvature of the curve, defined as
@@
Here prime denotes derivative with respect to $n$. Since we are only interested in the sign of $k$, it suffices to consider the function
@@
For $10<\eta<$ 1 this function is negative for all $n\geq 1$. This proves that the points $[R_{2}(n),R_{1}(n)]$ lie on a concave curve when $T=\frac{1}{2}$.

### v4

@@ VII Conclusions
In conclusion, we have shown that stellar ranks higher than $1$ can be certified with a very simple experimental setup that consists of a beam splitter, two click detectors, and an attenuator. The price to pay for the experimental simplicity is that the successful certification of higher stellar ranks is only possible for a limited set of states, in particular states that are sufficiently close to ideal Fock states. We have explained how the reduced total detection efficiency helps to certify stellar rank of higher Fock states, and we have numerically calculated the stellar rank thresholds for the considered stellar rank witnesses. For the simple witness based on probability that D1 clicks and D2 does not click we have observed that the maximum detectable stellar rank increases with decreasing detection efficiency and increases with increasing beam-splitter transmittance. The ability to certify higher stellar ranks can be further improved by considering more refined stellar rank witnesses that combine probability $R_{1}$ and probability $R_{2}$ that both detectors click simultaneously. In this latter case stellar ranks higher than $1$ can be certified even with perfect detectors and $\eta=1$ by utilizing an unbalanced scheme with $T\neq 0.5$. Nevertheless, even in such case losses can be helpful. Our results provide unique insights into the capabilities of the multiplexed click detectors and reveal that high stellar ranks can be certified even with a very simple scheme that involves only two detectors.
@@ Acknowledgements.
The author acknowledges support by the project OP JAC, Reg. No. CZ.02.01.01/00/23_021/0008790, of the Ministry of Education, Youth, and Sports of the Czech Republic and EU.
@@
Here we discuss the properties of a parametric curve specified by $R_{2}(n)$ and $R_{1}(n)$, where $n$ is now treated as a continuous parameter and $T=\frac{1}{2}$. Recall that
@@
which is positive for $0<\eta<1$. Since $R_{2}(n)$ monotonically increases with $n$, in order to prove the concavity it suffices to consider the signed curvature of the curve, defined as
@@
Here the prime denotes derivative with respect to $n$. Since we are only interested in the sign of $k$, it suffices to consider the function
@@
For $0<\eta<1$ this function is negative for all $n\geq 1$. This proves that the points $R_{2}(n),R_{1}(n)]$ lie on a concave curve when $T=\frac{1}{2}$.